Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $3.00 per unit. Enough capacity exists in the company’s plant to produce 30,800 units of the toy each month. Variable expenses to manufacture and sell one unit would be $1.90, and fixed expenses associated with the toy would total $50,320 per month. The company's Marketing Department predicts that demand for the new toy will exceed the 30,800 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $2,516 per month. Variable expenses in the rented facility would total $2.10 per unit, due to somewhat less efficient operations than in the main plant. Required:

1. What is the monthly break-even point for the new toy in unit sales and dollar sales.

2. How many units must be sold each month to attain a target profit of $11,520 per month?

3. If the sales manager receives a bonus of 20 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 26% return on the monthly investment in fixed expenses? (For all requirments, Round "per unit" to 2 decimal places, intermediate and final answers to the nearest whole number.)
Answer #1
Solution (1) monthly break-even point for the new toy in unit sales and dollar sales.
Existing Capacity Rented Total
Sales per unit 3 3
Variable cost -1.9 -2.1
Contribution per unit 1.1 0.9
Fixed cost per month 50320.00 2516.00
Contribution from existing capacity (30800*1.1) 33880
Less fixed expenses -50320.00
Net loss from existing capacity -16440
Beakeven
   =     Fixed cost of rent+loss from existing capacity
                     Contribution from new rented capacity
=(2516+16440)/0.9
= 21062.22 units
Hence break even sale = 30800+21062.22 = 51862.22
Solution (2 ) How many units must be sold each month to attain a target profit of $11,520 per month?
It is clear the profit can only generate profit when the company can sell in excess of 51861.22 units. Where existing capacity is 30800 units hence
Units required to be sold
   =Fixed cost of rent+loss from existing capacity+ target profit
                     Contribution from new rented capacity
Hence, Units required from new rented capacity
     =(16440+11520+2516)/0.9
    = 33862.22 units
Hence, companey need to sell   
     =(30800+33862.22)
     =64662.22
Proof
Existing Capacity Rented Total
Contribution per unit 1.1 0.9
Sale unit 30800 33862.22
Contribution ($) 33880 30476
Fixed cost -50320 -2516
Profit -16440 27960 11520
Solution (3) If the sales manager receives a bonus of 20 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 26% return on the monthly investment in fixed expenses?
Target profit = 26% of 2516
                         = 26%*2516
                         =654.16
Units from old capacity =30800 units
Loss from old capacity = 16440
Units to be sold from new capacity
    = (16440+2516+654.16)/0.9
    = 21789.07 units
Total units need to be sold = 30800+21789.07 =52589.07 units